The demand and supply functions of a commodity are given as follows:
Qd = 40 – lnp Qs = lnp – 20.
1.Find the equilibrium price and quantity
1
Expert's answer
2021-08-03T12:21:26-0400

For price to be at equilibrium ;"Q_d=Q_s"

"40-Inp =Inp - 20"


"2Inp=60"

"p=30"


"Quantity =40-30=10"